# Improving the Distributional Alignment of LLMs using Supervision

### v2

@@ 4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts and calibrating LLM-generated distributions with human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The alignment values for all models are shown in Appendix Table 7 for all the results in this section. All significance results are calculated via paired t-test with Bonferroni correction.
@@ 4.1 Does SD prompting improve alignment?
For RQ1, we compare SD vs. base prompts to study the effect of adding SD information on opinion alignment. Note that while prior work studies this question only for majority-voted responses Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024), we investigate the effect of adding SD information when comparing distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets.
@@ 4.2 Can we calibrate LLM distributions?
The results after regression, compared to pre-regression results, are shown in Table 1. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, improving by an average of 16.3%, where we average the delta between uncalibrated and calibrated opinion alignment across dataset-LLM-elicitation method settings. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment in these settings decreases by an average of 4.1%. These results suggest that LLM-generated distributions are somewhat uncalibrated, and a simple supervised regression can improve alignment.
Figure 2: Standard deviation vs. opinion alignment. Each point represents the average alignment for each dataset, LLM, and elicitation method. For visual clarity, we omit 43/290 uncalibrated points having opinion alignment below 60. Calibration tends to both increase opinion alignment and decrease standard deviation. It also decreases variance between settings.
Model Base prompt Sociodemographic prompt
$P$ $P_{cot}$ $P_{C}$ $S$ $S_{cot}$ $S_{C}$ $V$ $V_{cot}$ $V_{C}$ $P$ $P_{cot}$ $P_{C}$ $S$ $S_{cot}$ $S_{C}$ $V$ $V_{cot}$ $V_{C}$
@@
We also find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. Standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. Although there are some settings where uncalibrated distributions have a lower standard deviation than their calibrated counterparts, we emphasize that the variance between settings is much lower after calibration. While prior work Beck et al. (2024); Hu and Collier (2024) and our own results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results show us that calibration can reduce the variance between LLMs, datasets, and elicitation methods quite significantly. This means that with calibration, the choice of any particular model becomes less important.
We do not find distinct patterns with model size within model family (e.g., Mistral small vs. Mistral large). Indeed, smaller models are sometimes more aligned than larger models, though across prompting methods, elicitation methods, and datasets, larger models are more aligned more consistently (see Appendix Table 7). However, we do find that chain of thought prompting (Section 4.3) appears to increase alignment for smaller models, but not for larger models. The smaller Llama models in particular often refused to answer questions (i.e., questions would hit safety restrictions for attempting to role-play as a particular identity group), or would not follow instructions in formatting output distributions. We were thus unable to extract distributions for all questions, leading to higher variance in alignment. See Appendix Table 17 for percent distributions extracted in each setting.
@@
Table 3: Opinion alignment before ($V$) and after ($V_{C}$) calibration for three demographic categories (one from each dataset) using base-prompted, verbalized elicitation. “P.I” is political ideology and “I.” is income. Significantly higher values between each pair of columns (calculated via a paired t-test and Bonferroni correction) are bolded. The two “Average” columns on the right are averages across models, and the “All Demographics” rows are averages across the total set of SDs per dataset. See Appendix A.3 for all SDs. Calibrated distributions are more aligned with some SDs over others. Some models are better aligned with some datasets, e.g., Claude with OQA SDs and OLMo with WVS SDs.
4.3 The effect of chain of thought prompting
@@
5 How does alignment vary across SDs?
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it leads to the most aligned distributions with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 3. We show results for the same 5 LLMs used in Section 4. See Appendix A.3 for alignment for other demographics.
5.1 The effect of calibration
First, we compare opinion alignment between base and calibrated distributions. We use the regression models trained on all SD groups and study how such aggregate regression models might calibrate individual sociodemographics. We note few significant differences (calculated with a paired t-test and Bonferroni correction) as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more absolutely aligned than their uncalibrated counterparts for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. This includes all demographic categories, including those in Table 3 and Appendix A.3. The delta improvement across demographics is most significant for OLMo on the WGM and OQA datasets. As expected, calibration improves alignment for some demographics and reduces alignment for others. We find that it tends to improve alignment of SD groups that were less aligned with base-prompted distributions, such as Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment decreases for very high aligned demographic groups, such as Aus/NZ and Northern Europe.
@@ 5.1 The effect of calibration
5.2 Alignment variation across LLMs
We now compare opinion alignment between different LLMs. For world region, the most aligned model is dependent on region, though Claude-generated distributions are most aligned with Africa and South/Central America, while Mistral and Llama are most aligned with Europe, Asia, and the Middle East. Qwen is most aligned with North America and Western Europe before calibration, despite being more extensively pretrained on Chinese data in addition to English data. For political ideology, Claude appears to be most aligned with all ideologies, followed by Qwen. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. This is discussed further in Section 7.
In general across sociodemographics, the Claude-generated distributions tend to be more highly aligned with the OQA dataset over the others. All models have lowest alignment with middle income populations. This might be because low and high income populations have opinions on these survey questions that are less varied than people with middle income. In other words, income might be less important in determining survey answers for those with middle income.
6 Does supervision work with less data?
6.1 Amount of supervised data required
As discussed in Section 3.6, we evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full supervised examples for each dataset, and average over 10 random samples for each size. We find that as few as 1-10 examples (with $\approx$4 answer choices per example) can suffice to closely match the minimal MSE for any particular dataset-LLM-elicitation setting. Figure 3, plotting Mean Squared Error (MSE) over training data size, shows that MSE usually converges between 1 and 10 examples, though this is model and dataset dependent.
6.2 Comparison with few-shot prompting
Given that the regression models appear to converge with so few examples, one might ask whether the regression is actually learning anything useful, or if we could achieve a similar improvement with few shot prompting instead. To investigate this, we 5-shot prompt the LLMs with the same examples given to the regression model in the 5 example calibration setting. We use a 5-shot prompt since this lies in the middle of how many examples were needed for convergence for the regression models. Whereas we averaged over 10 random samples in those experiments, here we use only 3 of the random samples because we found little variation. We average alignment over the 3 runs.
6.3 The effect on individual SD groups
Although we see convergence with fewer examples, does using a small set of random examples affect alignment of individual demographics? We calibrate with a set of 5 examples (in the middle of the 1-10 range described above for MSE convergence). Although degradation on average is close to zero, individual demographics are affected differently. Table 5 shows the five SDs most affected and the degradation amount for both base and SD-prompted verbalized distributions. We see that SD prompting affects which demographics have the highest difference in alignment between the full amount of data and with a random sample of 5 examples, although none of the differences are statistically significant. Using SD information brings more degradation for most affected demographics but changes the most affected SDs from those that are historically less represented (SE Asia, less than high school education, Black, and Hispanic SD groups) to those that are more highly represented (Europe, Aus/NZ, tertiary education SD groups).
7 Does post-training impact alignment?
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + supervised finetuning (SFT), base + SFT + direct preference optimization (DPO), and the instruct model (base + SFT + DPO + reinforcement learning with verifiable rewards (RLVR)). The SFT step involves finetuning on instruction datasets and scaled synthetic data, the DPO step involves preference finetuning with DPO, and the RLVR step involves finetuning targeted at domains where prompts with verifiable rewards can be constructed, such as math OLMo et al. (2025). We use these four models out-of-the-box via Huggingface and study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it is typically most aligned.
Category Demographic $V_{F}$ $V_{5}$ $\Delta$
Base Region Southeast Asia 81.0 79.4 1.6
Education Less than HS 83.1 81.6 1.5
Race Black 86.6 85.1 1.5
Race Hispanic 87.2 85.9 1.4
Pol. Party Democrat 84.7 83.3 1.4
SD Region Northern Europe 87.1 84.6 2.4
Region Western Europe 87.1 84.9 2.3
Region Aus/NZ 84.3 82.1 2.2
Education Tertiary 88.6 86.5 2.1
Employment Unemployed 86.9 84.9 2.0
8 Discussion
Limitations of demographic modeling. Modeling groups wrt. their demographics assuredly provides only a narrow, incomplete view of people, and demographics alone do not determine our opinions. That said, such models may usefully approximate distributions of group opinions, provided care is exercised. We note that our methods for improving alignment should not be viewed as a way to replace human respondents in surveys, but rather a step towards more accurately representing groups.
The effect of SD prompting. We showed that simply prompting with SD information does not necessarily lead to more aligned distributions. Future work might study eliciting distributions with more implicit demographic information or with past opinions as prior work has done with single responses Hwang et al. (2023); Do et al. (2025). One could make personalized predictions for a sample of a group’s individuals members Gordon et al. (2022); Hwang et al. (2023) rather than a single distributive prediction for the group. Multimodal prompts might also add SD information implicitly, though this might come with additional biases.
The choice of LLM and distribution elicitation method matters less with calibration. We find that distributional opinion alignment is dataset, LLM, and distribution elicitation method dependent, consistent with prior work evaluating majority responses Beck et al. (2024). However, verbally-elicited distributions are most aligned with human distributions in the most settings. We also find that the Claude 3.5 models are most aligned with human responses on average across all datasets, though we highlight that this is also dataset-dependent. More importantly however, we find that calibration reduces variance in alignment across models and methods; i.e., the choice of model and elicitation method matters less after calibration.
9 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques, as well as increase consistency between settings. Our benchmark also serves to help enable and stimulate future research.
10 Limitations
Prompting methods. We do not experiment with varied temperatures for our generations. For elicitation methods, though a larger $n$ might improve generated distributions, our choice of $n$ is reasonable given inference costs and following prior work Xiong et al. (2024). We use $n=3$ for verbalized elicitation (as opposed to $n=5$) since we found little variance between runs.
Logit-based elicitation methods. We focus on methods that broadly work with open-source, open-model, and API-access LLMs. That said, we report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix A.2). We found that log probabilities with calibration yielded the most aligned distributions in 10/24 settings considered. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied. We only studied a subset of the demographics available in the datasets. Future work might study additional and/or intersectional demographics Crenshaw (1991).
Regression. Our regression predicts each answer choice individually and then normalizes the predicted answers. We tried constraining optimization to learn weights for all answer choices simultaneously and enforce proper distributions, but learning the answer choices individually performed better. We also optimize regression for MSE instead of the alignment metric. Future work might also ensemble multiple elicitation methods.
Ethical Considerations
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for creating LLMs that more accurately portray and reason about population groups, this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation.
We note that by prompting LLMs with SD information, and evaluating their distributions against humans with those SDs, we are essentializing the SD to the identities of the survey respondents and their opinions Wang et al. (2025a). Although SDs are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). By seeking to infer distributions of human responses rather than majority responses for each SD group, we seek to align LLMs with groups in a distributionally pluralistic manner, rather than assume all members of the group would answer similarly.
Acknowledgments
Appendix A Additional Results
A.1 Alignment results for all models

### v3

@@ 4 Aligning LLMs with human opinions
We compare two methods: adding SD information to prompts vs. calibrating LLM-generated distributions with human ground truth data for supervision. We evaluate the effectiveness of these two methods in aligning with opinion distributions of those SD groups across models, datasets, and probability elicitation methods. Our main results are shown in Table 1. The alignment values for all models are shown in Appendix Table 7 for all the results in this section. All significance results are calculated via paired t-test with Bonferroni correction.
@@ 4.1 Does SD prompting improve alignment?
For RQ1, we compare SD vs. base (SD-agnostic) prompts to study the effect of adding SD information on opinion alignment. While prior work studies this question only for majority-voted responses Hu and Collier (2024); Sun et al. (2025a); Mukherjee et al. (2024); Masoud et al. (2025); AlKhamissi et al. (2024), we investigate the effect of adding SD information when comparing distributions. We study whether this is consistent across distribution elicitation methods, models, and datasets.
@@ 4.2 Can we calibrate LLM distributions?
The results after regression, compared to pre-regression results, are shown in Table 1. Calibration increases opinion alignment in 94.8% of dataset-LLM-elicitation method settings, improving by an average of 16.3%, where we average the delta between uncalibrated and calibrated opinion alignment across dataset-LLM-elicitation method settings. We find that in the 5.2% of settings where it does not increase opinion alignment, the original aligment is relatively high and the decrease in alignment with calibration is low: alignment in these settings decreases by an average of 4.1%. These results suggest that LLM-generated distributions are somewhat uncalibrated, and simple supervised regression can improve alignment.
Figure 2: Standard deviation vs. opinion alignment. Each point represents the average alignment for each dataset, LLM, and elicitation method setting. For visual clarity, we omit 43/270 uncalibrated points having opinion alignment below 60. Calibration decreases variance in opinion alignment and standard deviation between settings.
@@
We also find that calibration reduces variance across settings. Figure 2 plots standard deviation vs. opinion alignment, showing standard deviation across datasets, LLMs and elicitation methods before and after calibration. Standard deviation is over 3 times smaller across all settings, and up to 5 times smaller per dataset. Although there are some settings where uncalibrated distributions have a lower standard deviation than their calibrated counterparts, we emphasize that the variance between settings is much lower after calibration. While prior work Beck et al. (2024); Hu and Collier (2024) and our own results in Section 4.1 showed that opinion alignment is dataset, LLM, and elicitation method dependent, these results indicate that calibration can improve the consistency between LLMs, datasets, and elicitation methods quite significantly. Differences between models become less pronounced following calibration.
We do not find distinct patterns with model size within model family (e.g., Mistral small vs. Mistral large). Indeed, smaller models are sometimes more aligned than larger models, though across prompting methods, elicitation methods, and datasets, larger models are more aligned more consistently (see Appendix Table 7). However, we do find that chain of thought prompting (Appendix B) appears to increase alignment for smaller models, but not for larger models. The smaller Llama models in particular often refused to answer questions (i.e., questions would hit safety restrictions for attempting to role-play as a particular identity group), or would not follow instructions in formatting output distributions. We were thus unable to extract distributions for all questions, leading to higher variance in alignment. See Appendix Table 13 for percent distributions extracted in each setting.
@@
Table 2: Opinion alignment before ($V$) and after ($V_{C}$) calibration for three demographic categories (one from each dataset) using base-prompted, verbalized elicitation. “P.I” is political ideology and “I.” is income. Significantly higher values between each pair of columns (calculated via a paired t-test and Bonferroni correction) are bolded. The two “Average” columns on the right are averages across models, and the “All Demographics” rows are averages across the total set of SDs per dataset. See Appendix D.3 for all SDs. Calibrated distributions are more aligned with some SDs over others. Some models are better aligned with some datasets, e.g., Claude with OQA SDs and OLMo with WVS SDs.
5 How does alignment vary across SDs?
We focus on three demographic categories, one from each dataset: world region (WGM), political ideology (OQA), and income (WVS). We only look at distributions elicited with the verbalized method, as it leads to the most aligned distributions with most models across datasets. We use the base-prompted distributions, as we find SD-prompted distributions do not show consistent improvement over base-prompted distributions for individual sociodemographics. Opinion alignment for world region (WGM), political ideology (OQA), and income (WVS) is shown in Table 2. We show results for the same 5 LLMs used in Section 4. See Appendix D.3 for alignment for other demographics.
5.1 The effect of calibration
We use the regression models trained on all SD groups to study how such aggregate models might calibrate individual SDs. We note few significant differences as our sample size for each demographic is low (4-6 examples). However, calibrated distributions are more absolutely aligned than their uncalibrated counterparts for 73.57% of demographics in WGM, 68.41% in OQA, and 78.17% in WVS, across models for base verbalized distributions. This includes all demographic categories, including those in Table 2 and Appendix D.3. The delta improvement across demographics is most significant for OLMo on the WGM and OQA datasets.
As expected, calibration improves alignment for some demographics and reduces alignment for others. We find that it tends to improve alignment of SD groups that were less aligned with base-prompted distributions, such as Central America/Mexico, South America, Southern Africa, and Western Africa for world region. Alignment decreases for very high aligned demographic groups, such as Aus/NZ and Northern Europe.
5.2 Alignment variation across LLMs
For world region, alignment varies by model: Claude is most aligned with Africa and South/Central America, while Mistral and Llama are more aligned with Europe, Asia, and the Middle East. Qwen is most aligned with North America and Western Europe pre-calibration despite more extensive Chinese pretraining. For political ideology, Claude is most aligned with all ideologies, followed by Qwen. On income (and most demographics in the WVS dataset), OLMo is most aligned with all levels, surprising given its lower alignment on the demographics in other datasets. See Appendix A for further discussion of OLMo alignment and post-training effects.
Across sociodemographics, Claude-generated distributions tend to be more highly aligned with the OQA dataset. All models show lowest alignment with middle income populations, possibly because low and high income groups have less varied opinions on these questions, i.e., income may be less predictive for middle income respondents.
6 Does supervision work with less data?
As discussed in Section 3.6, we evaluate calibrated opinion alignment with 1, 5, 10, 50, 100, 200, …, full examples for each dataset, and average over 10 random samples for each size. We find that as few as 1-10 full examples can suffice to closely match the minimal Mean Squared Error (MSE) for any particular dataset-LLM-elicitation setting. Plotting MSE over training data size shows that MSE usually converges between 1-10 full examples, though this is model and dataset dependent (see Appendix F). We find that in comparison with few-shot prompting (Appendix C), alignment improvement is more consistent with calibration.
6.1 The effect on individual SD groups
Although we see convergence with fewer examples, does using a small set of random examples affect alignment of individual demographics? We calibrate with a set of 5 full examples (in the middle of the 1-10 range described above for MSE convergence). While average degradation is near zero, individual demographics are affected differently.
Table 3 shows the five most effected SDs and degradation for base and SD-prompted distributions. SD prompting changes which SDs show the highest alignment differences between full data and 5-full-example samples, though no differences are statistically significant. Using SD information increases degradation for the most affected demographics but shifts impact from historically underrepresented groups (SE Asia, less than high school education, Black, Hispanic) to overrepresented ones (Europe, Aus/NZ, tertiary education).
Category Demographic $V_{F}$ $V_{5}$ $\Delta$
Base Region Southeast Asia 81.0 79.4 1.6
Education Less than HS 83.1 81.6 1.5
Race Black 86.6 85.1 1.5
Race Hispanic 87.2 85.9 1.4
Pol. Party Democrat 84.7 83.3 1.4
SD Region Northern Europe 87.1 84.6 2.4
Region Western Europe 87.1 84.9 2.3
Region Aus/NZ 84.3 82.1 2.2
Education Tertiary 88.6 86.5 2.1
Employment Unemployed 86.9 84.9 2.0
7 Discussion
The effect of SD prompting. Given much prior work with conflicting findings, our extensive experiments show convincingly that prompting with SD information alone fails to consistently yield more aligned distributions. Future work might explore eliciting distributions with implicit demographic information or past opinions Hwang et al. (2023); Do et al. (2025), personalization to individual group members rather than entire groups Gordon et al. (2022); Hwang et al. (2023), or with multimodal prompts to incorporate implicit information.
The choice of LLM and distribution elicitation method matters less with calibration. We find that distributional alignment is dataset, LLM, and elicitation method dependent, consistent with prior work evaluating majority responses Beck et al. (2024). However, verbally-elicited distributions are most aligned with human distributions in the most settings. In general, we find that calibration improves alignment more reliably across models and methods; i.e., the choice of model and elicitation method matters less after calibration.
8 Conclusion
We investigate LLM distribution alignment with human responses to subjective large-scale surveys, both on average and across diverse population groups. We show that using simple supervision can improve alignment with population groups consistently across datasets, models, and distribution elicitation techniques, as well as increase consistency between settings. Our benchmark serves to help enable and stimulate future research.
Limitations
Demographic modeling
Modeling groups wrt. their demographics assuredly provides only a narrow, incomplete view of people, and demographics alone do not determine our opinions. That said, such models may usefully approximate distributions of group opinions, provided care is exercised. We note that our methods for improving alignment should not be viewed as a way to replace human respondents in surveys, but rather a step towards more accurately representing groups.
Calibration model
Our calibration model predicts each answer choice individually and then normalizes the predicted answers, ignoring any correlations among answer choices. This might introduce risks of over-smoothing, loss of distributional structure, and potential distortions. Interestingly, though we tried constrained optimization that enforced proper distributions across all choices simultaneously, learning answer choices individually performed better. Future work might further analyze the effects of this individual answer choice prediction on alignment.
Prompting methods
We only use a single temperature setting for our generations; future work might further analyze the effects of temperature on the quality of LLM-generated distributions. For elicitation methods, though a larger $n$ might improve generated distributions, our choice of $n$ is reasonable given inference costs and following prior work Xiong et al. (2024). We use $n=3$ for verbalized elicitation (as opposed to $n=5$) since we found little variance between runs.
Logit-based elicitation methods
We focus on methods that broadly work with open-source, open-model, and API-access LLMs. That said, we report a smaller scale study with logit-based distribution elicitation with four Llama models (Appendix D.2). We found that log probabilities with calibration yielded the most aligned distributions in 10/24 settings considered. Future work might look more into comparisons between logit-based and verbalized distribution elicitation.
Demographics studied
We only studied a subset of the demographics available in the three datasets we use due to cost limitations. Future work might study additional and/or intersectional demographics Crenshaw (1991).
Ethical Considerations
In this work, we study how we might align LLMs to the survey response distributions of various social groups. While this can be valuable for creating LLMs that more accurately portray and reason about population groups, this can also be used for adversarial purposes, such as chat bots attempting to simulate different user groups, targeted advertising, or targeted misinformation. We do not intend our datasets, models, and methods to be used for these adversarial purposes. Even for the intended purpose of furthering research in alignment of LLMs to population groups, we emphasize careful consideration of use-case and using caution to avoid propagating bias and stereotypes.
We note that by prompting LLMs with SD information, and evaluating their distributions against humans with those SDs, we are essentializing the SD to the identities of the survey respondents and their opinions Wang et al. (2025). Although SDs are not at all the only factor that influences people’s opinions, they certainly play a part Sap et al. (2022); Biester et al. (2022); Pei and Jurgens (2023). By seeking to infer distributions of human responses rather than majority responses for each SD group, we seek to align LLMs with groups in a distributionally pluralistic manner, rather than assume all members of the group would answer similarly.
Acknowledgments
Appendix A Impact of post-training on alignment
To study effects of post-training methods for RQ5, we compare differences in performance between four OLMo-2-7B models (the only completely open source model family we evaluate): the base model, base + supervised finetuning (SFT), base + SFT + direct preference optimization (DPO), and the instruct model (base + SFT + DPO + reinforcement learning with verifiable rewards (RLVR)). The SFT step involves finetuning on instruction datasets and scaled synthetic data, the DPO step involves preference finetuning with DPO, and the RLVR step involves finetuning targeted at domains where prompts with verifiable rewards can be constructed, such as math OLMo et al. (2025). We use these four models out-of-the-box via Huggingface  [^3] https://huggingface.co/allenai/models and study differences between base vs. SD and calibrated vs. uncalibrated with the verbalized elicitation method, since it is typically most aligned.
Appendix B The effect of chain of thought prompting
@@
Model Base prompt Sociodemographic prompt
$P$ $P_{cot}$ $P_{C}$ $S$ $S_{cot}$ $S_{C}$ $V$ $V_{cot}$ $V_{C}$ $P$ $P_{cot}$ $P_{C}$ $S$ $S_{cot}$ $S_{C}$ $V$ $V_{cot}$ $V_{C}$
Appendix C Comparison with few-shot prompting
Given that the regression models appear to converge with so few examples, one might ask whether the regression is actually learning anything useful, or if we could achieve a similar improvement with few shot prompting. To investigate this, we 5-shot prompt the LLMs with the same examples given to the regression model in the 5 full example calibration setting. We use a 5-shot prompt since this lies in the middle of the number of full examples that were needed for convergence for the regression models. Whereas we averaged over 10 random samples in those experiments, here we use only 3 of the random samples because we found little variation. We average alignment over the 3 runs.
@@ Appendix C Comparison with few-shot prompting
Appendix D Additional Results
D.1 Alignment results for all models